OpenStudy (anonymous):

Find the equation of the line that is the perpendicular bisector of the segment that connects the points (3,6) and (-2,4)

OpenStudy (anonymous):

\[(y^2 - y^1) \over (x^2 - x^1)\] to find slope y-y1=m(x-x1) to find equation of the line Flip numerator/denominator and sign to find perpendicular lines

OpenStudy (anonymous):

perpendicular bisector of the segment is nerd talk for "find the perpendicular line" im assuming

OpenStudy (anonymous):

slope= 2/5 y-6=2/5 (x-3) y=2/5x +24/5 ? Is that all i have to do?

OpenStudy (anonymous):

that is the line that runs through the points given in the problem, they want to you find one that runs perpendicular to it

OpenStudy (anonymous):

I cant remember, which part do i flip?

OpenStudy (anonymous):

m, or the slope

OpenStudy (anonymous):

so -5/2 is the new slope

OpenStudy (anonymous):

ok thx